OpenStudy (anonymous):

Determine whethter the equation is linear in x. 2√x+5=10

OpenStudy (goformit100):

"Welcome to OpenStudy. I can guide you. Please use the chat for off topic questions. And remember to give the helper a medal, by clicking on "Best Answer". We follow a code of conduct, ( http://openstudy.com/code-of-conduct ). Please take a moment to read it."

OpenStudy (jmark):

Given 2√x+5=10 2√x=10-5 √x=5/2 x=25/4 then equation become 2*(5/2)+5=10 10=10 there for the equation is linear in x